OpenStudy (purple_pink):

find 2 ratios that are equivalent to the given ratio 50 : 75

OpenStudy (purple_pink):

@TheProfessor35

OpenStudy (purple_pink):

@maddisonGruby

OpenStudy (theprofessor35):

Alright, lets start out by picking any number, lets pick 2. 50*2 = ? 75*2 = ?

OpenStudy (theprofessor35):

100/150 , since we multiplied

OpenStudy (theprofessor35):

Now, lets find one more. I choose 3. 50*3 = ? 75*3 = ?

OpenStudy (theprofessor35):

150/225

OpenStudy (theprofessor35):

Does that clear it up? We can also divide!

OpenStudy (theprofessor35):

50/5 = 10 75/5 = 15 10:15